Mt. Pinatubo, in the Philippines, erupted in 1991. Sulfure was abundant after the eruption. Sulfur is formed from the mixing of two common volcanic gases, SO2, H2S with liquid water and S8(s) as products. If 4.5e5 Kg of sulfure is formed during the eruption and the pressure was 730.36 Torr and 295 K what was the minimum volume of each gas that must have been available during the eruption?
1
Expert's answer
2016-02-24T00:01:01-0500
16H2S + 8SO2 = 3S8 + 16H2O
N(S8)=4.5e8/256=1.758e3 moles
N(SO2)=1.758e3*8/3= 4.688e3 moles
N(H2S)=1.758e3*16/3= 9.375e3 moles
PV=(m/M)RT V=(NRT)/P
730.36 Torr= 97,373 KPa
V(SO2)=4.688e3*8,314*295/97,373=1,18e5 L
V(H2S)=9.375e3*8,314*295/97,373=2,36e5 L